OpenStudy (anonymous):

what is (3c^2d^5)(2c^3d^0)^2 in simplified form

OpenStudy (anonymous):

well its multiple choice and my choices are a.6c^8d^7 b.12c^7d^5 c.12c^8d^5 and d.6c^12d

OpenStudy (unklerhaukus):

\[(3c^2d^5)\times(2c^3d^0)^2\]\[=3c^2d^5\times4c^6\]\[=12c^8d^5\]\[=c.\]

OpenStudy (anonymous):

thanks so much for the help...